OpenStudy (anonymous):

Help Please When graphing the following equations, which one will be a straight line? A. y = x2 + 2 B. f(x) = x + 1 C. y = |x + 4| D. x2 + y2 = 1

OpenStudy (mikurout):

option b and c.

OpenStudy (anonymous):

it would be B right?

OpenStudy (mikurout):

i think in option a x2 = x^2 and in option d x2 = x^2

OpenStudy (anonymous):

wait im lost

OpenStudy (mikurout):

why not option c?

OpenStudy (imstuck):

not c because it's an absolute value function and they have the general shape of this:only I believe this one is moved up 4 units.

OpenStudy (mikurout):

use the general equation y=mx +c

OpenStudy (imstuck):

A is a parabola. Not a straight line.

OpenStudy (imstuck):

D is a circle. Not a straight line

OpenStudy (imstuck):

The only straight line is B.

OpenStudy (mikurout):

@IMStuck yes u are right. I was thinking just straight lines. means which are not curves.

OpenStudy (imstuck):

TY for the medal!

OpenStudy (mikurout):

welcome.